A valószínűségszámítási gyakorlaton megajánlott jegyek:

diákok:

Andrékó Judit  2
Bácskai Marianna  3
Balázs Erika  3
Bálint Melinda  4
Balogh Péter  −
Benczik Attila  −
Boros Ildikó  4
Czibere Marianna  2
Császár Katalin  3
Dienes Zoltán  −
Éliás Anikó  4
Fórizs Fruzsina  2
Föczén Beáta  3
Gőgh Timea  2
Gombkötő Angéla  −
Holló Anikó  −
Huszty Árpád  −
Iván Éva  −
Kelemen Imre  3
Kerékgyártó Zita  3
Kocsis Zsuzsanna  3
Koncz Nikoletta  3
Lovas Rezső László  5
Lukács Erika  2
Orosz Erzsébet  2
Pelei Csaba  ?
Pintér Györgyi  2
Szabó Erika  4
Szilágyi Renáta Judit −
Sztankó Paticia  3
Szűcs Mónika  2
Turák Olivér  2
Villás Rita  3

Major Péter